is a coupon bond with no maturity. The price of a perpetuity (in dollars) t it pays an annual coupon of $90 and its yield to maturity is 1 1% is S 818 18 response to the nearest two decimal,place) (Roundy r The yield to maturity on a consol that pays an annual coupon of $50 and sells for $400 is%.(Round your response to the nearest whole number)
Answer #1

1. A consol is a coupon bond with no maturity.

2. Ans: $818.18

Explanation:

P = A / i

90 = 90 / 0.11 = $818.18

3. Ans: 12.5%

Explanation:

i = (A / P) * 100

= (50 / 400) * 100

= 12.5%